How long would it take a car traveling at a speed of 40 mi/h to reach a town 160 miles away?
Found 2 solutions by Theo, dkppathak:
Answer by Theo(13342)

rate * time = distance.
rate = 40 miles per hour.
distance = 160 miles
rate * time = distance becomes 40 * time = 160.
solve for time to get time = 160 / 40 = 4 hours.

Answer by dkppathak(439)

How long would it take a car traveling at a speed of 40 mi/h to reach a town 160 miles away?
solution
speed=distance/time
or time=distance /speed
time =160/40=4 hours